# On the Identification of Diagnostic Expectations: Econometric Insights from DSGE Models

### v1

@@ Abstract
This paper provides the first econometric evidence for diagnostic expectations (DE) in DSGE models. Using the identification framework of Qu and Tkachenko (2017), I show that DE generate dynamics unattainable under rational expectations (RE), with no RE parameterization capable of matching the volatility and persistence patterns implied by DE. Consequently, DE are not observationally equivalent to RE and constitute an endogenous source of macroeconomic fluctuations, distinct from both structural frictions and exogenous shocks. From an econometric perspective, DE preserve overall model identification but weaken the identification of shock variances. To ensure robust conclusions across estimation methods and equilibrium conditions, I extend Bayesian estimation with Sequential Monte Carlo sampling to the indeterminacy domain. These findings advance the econometric study of expectations and highlight the macroeconomic relevance of diagnostic beliefs.
JEL classification: C11; C13; C54; C63; E71
Keywords: Diagnostic expectations; DSGE; Identification;
@@ 1 Introduction
I begin with a small-scale DSGE framework and show that the diagnostic parameter can be identified globally using the frequency domain methodology of Qu and Tkachenko (2017). This approach introduces a frequency domain expression for the Kullback–Leibler (KL) distance between two DSGE models, which measures the difference between their implied spectra (the frequency domain representation of the autocovariance functions). If no two parameter sets generate identical spectrum, that is, if the KL distance is strictly positive, the model is globally identified. Applying this criterion, I find that incorporating DE does not compromise overall model identification but systematically weakens the identification of shock variances. By contrast, key structural parameters remain well identified. This pattern is consistent with the theoretical structure of DE: the diagnostic parameter alters only the shock impact coefficients while leaving autoregressive dynamics unchanged, thereby reducing the sensitivity of the model implied spectrum to changes in shock variances.
@@
Related literature.   This paper relates to two strands of literature: the literature on DE and the broader research on the identification of structural parameters in macroeconomic models. The first strand relates to the growing body of research on DE. Bordalo et al. (2018) (BGS) introduce DE to explain several features of credit cycles, demonstrating how agents’ psychological tendency to overweight representative future outcomes amplifies economic fluctuations. Bordalo et al. (2020) explore DE with dispersed information, finding that agents overreact to private signals while underreacting to consensus forecasts. Bordalo et al. (2021) incorporate price learning and speculative behavior, accounting for the underreaction-overshooting-crash pattern in asset price bubbles. Guo et al. (2023) examine DE under incomplete information and document excess consumption sensitivity consistent with survey evidence. More recently, Bianchi et al. (2024a) show that a DE-based RBC model better replicates boom-bust cycles compared to its rational expectations counterpart. Na and Yoo (2025) apply this approach to study countercyclical external balances in emerging markets. L’Huillier et al. (2024) incorporate DE into a New Keynesian framework, providing a foundation for my identification analysis. While recent contributions propose smooth DE that distortion vary with uncertainty (Bianchi et al., 2024b), this paper employs the standard BGS framework, which remains the theoretical foundation for most research in this area.
@@ 3.2.4 Global identification
Under DE, the shock variance parameters $\sigma_{a}$ and $\sigma_{g}$ exhibit the weakest identification strength, followed by $\phi_{\pi}$. As shown in L’Huillier et al. (2024)’s online appendix [^9] See also the analytical solution to the modified model in Appendix section 6.1., the DE parameter $\theta$ influences the policy functions solely through the shock terms. Consequently, introducing $\theta$ effectively redistributes the explanatory power of shocks on endogenous variables, which reduces the curvature of the model implied spectrum to changes in shock variances. Intuitively, since $\theta$ modulates how shocks propagate into the system, variations in the underlying shock variances produce smaller changes in the observable spectral properties. This reduction in spectral sensitivity translates into weaker identification for $\sigma_{a}$ and $\sigma_{g}$. Despite this, the model remains globally identifiable under DE, with the DE parameter $\theta$ itself demonstrating relatively strong identification.
@@ 3.3 Identification under indeterminacy
Notes: The inverse Gamma prior is given by $p(\sigma\mid\iota,s)\propto\sigma^{-\iota-1}e^{-\frac{s^{2}}{2\sigma^{2}}}$, where $\iota=4$ and $s=0.75$. Posterior results were estimated using a modified MATLAB code from Herbst and Schorfheide (2016). All settings are consistent with the determinacy settings in Dynare: number of particles ($N$) = 3,000, number of stages ($N_{\phi}$) = 200, $\lambda=2$, initial scaling parameter = 0.5, and initial acceptance rate = 0.25. Mutation block = 1. The modifications extend the algorithm to cover the indeterminacy region, manage random seeds deterministically across parallel workers to ensure fully reproducible results, and implement numerically stable weight updating via the log-sum-exp transformation to prevent overflow. The estimation results are robust to increasing the number of stages or particles, as well as to employing diffuse priors for the projection coefficient.
@@ 4.1 Global identification
To estimate the model, L’Huillier et al. (2024) use 12 observables: output growth, consumption growth, investment growth, inflation, the interest rate, wage growth, labor growth, and one-period-ahead forecasts of output growth, consumption growth, investment growth, inflation, and the interest rate. However, as noted by Qu and Tkachenko (2023), global identification analysis requires the spectral density to be non-singular. Following Qu and Tkachenko (2023), I therefore work with log deviations from steady states of output, consumption, investment, wage, and labor, rather than their growth rates. Together with inflation and the interest rate, this yields 7 measurement variables in total. For completeness, Appendix 6.2 contains a detailed discussion of the modified information friction block as well as the full set of model equations.
Using the same approach as in the small-scale DSGE analysis, I take the posterior mean from Bayesian estimation as the benchmark parameter vector, denoted $\gamma_{0}^{med}$. Since the information-friction block is modified, I re-estimate the medium-scale model using standard MCMC sampling. [^11] The resulting posterior distribution is very similar to that reported in Table 1 of L’Huillier et al. (2024), with the only notable difference being the variance of the noise shock, an expected outcome given that the signal now pertains to the growth rate rather than the level of permanent productivity. The full posterior results are reported in Table 14.
@@ 4.2 Observable Equivalence and the Role of Frictions
First, I examine observable equivalence between the DE and RE specifications. The KL distance is 0.99 for $T=80$ and 1 for $T=150$, indicating that no observable equivalence exists between DE and RE, even with a rich set of frictions. Among the parameters that shift most significantly, the medium-scale RE model requires weaker habit formation, more flexible prices and wages, more volatile markup shocks, and a more dovish monetary policy response to inflation in order to approximate the spectrum of the DE model. In addition, the inverse Frisch elasticity decreases substantially (from 3.71 to 2.39), implying a much more elastic labor supply under RE. These adjustments are intuitive. To replicate the stronger demand responses induced by DE consumers’ extrapolative expectations, RE consumers must rely less on past consumption habits. Since the RE Phillips curves lack extrapolative dynamics, prices must adjust more flexibly than under DE to reproduce observed behavior. Likewise, markup shocks become more volatile in the RE case because, without endogenous amplification from expectations, the model depends more heavily on exogenous shock variances. The sharp increase in labor supply elasticity reflects the RE model’s need for structural amplification to replicate the employment volatility that DE generates through behavioral amplification. Since RE agents do not overreact to shocks, workers must be far more responsive to wage changes to achieve comparable employment fluctuations. Finally, as argued by L’Huillier et al. (2024), when the intertemporal elasticity of substitution equals one, a hawkish monetary policy rule dampens the propagation of noise shocks in RE models. Hence, to mimic DE dynamics, the RE model requires an unrealistically dovish policy rule.
@@
To further study the role of information frictions, I fix $\sigma_{s}$ at alternative levels and allow the remaining parameters of the DE model to adjust in order to fit the benchmark DE spectrum. In particular, I compare a near-frictionless case ($\sigma_{s}=0.01$) with a high-friction case ($\sigma_{s}=1$). This exercise highlights how informational frictions reshape both the structural parameters and the degree of DE distortion required to replicate the benchmark dynamics. For the near-frictionless case, no observable equivalence is detected: the KL distance is 0.50 for $T=80$ and 0.719 for $T=150$. The parameters remain close to those of the benchmark DE model. In particular, the DE distortion parameter decreases only slightly, from 0.73 to 0.70. The more notable changes occur in investment-related parameters. When DE agents can observe productivity growth almost perfectly, investment decisions respond less to spurious signals. As a result, the model requires less volatility from marginal efficiency shocks. For the high-friction case, no observable equivalence is detected either: the KL distance equals 1 for both $T=80$ and $T=150$. With severe information frictions, the underreaction channel becomes stronger, which in turn raises the required degree of diagnostic distortion. In this setting, the model relies more heavily on volatile investment specific shocks to generate sufficient fluctuations, while at the same time making capital utilization efficiency more costly to adjust. These adjustments reflect the economy’s need for stronger structural amplification when informational noise is large.
@@
$\gamma_{0}^{med}$ $\theta=0$ $\sigma_{s}=0.01$ $\sigma_{s}=1$ $\kappa_{p}=0.1$ $\kappa_{w}=0.1$ $h=0.1$
KL 0.1160 0.0066 0.0481 0.0743 0.1060 0.6520
$T=80$ 0.9940 0.5000 1.0000 0.9600 0.9910 1.0000
$T=150$ 1.0000 0.7190 1.0000 0.9980 1.0000 1.0000
$\theta$ 0.73 0.00 0.70 0.92 0.36 0.46 0.10
@@
$\frac{\chi^{\prime\prime}(1)}{\chi^{\prime}(1)}$ 5.09 4.94 4.48 6.60 4.99 3.96 4.98
@@
In sum, these experiments show that DE generates dynamics distinct from RE. While other frictions can partially substitute for DE distortions by shifting extrapolation to other channels, none can fully replicate the benchmark DE dynamics. This suggests that DE captures a distinct endogenous amplification mechanism that cannot be mimicked by conventional frictions alone, underscoring its importance in explaining macroeconomic fluctuations.

### v2

@@ Abstract
This paper provides the first econometric evidence for diagnostic expectations (DE) in DSGE models. Using the identification framework of Qu and Tkachenko (2017), I show that DE generate dynamics unreplicable under rational expectations (RE), with no RE parameterization capable of matching the autocovariance implied by DE. Consequently, DE are not observationally equivalent to RE and constitute an endogenous source of macroeconomic fluctuations, distinct from both structural frictions and exogenous shocks. From an econometric perspective, DE preserve overall model identification but weaken the identification of shock variances. To ensure robust conclusions across estimation methods and equilibrium conditions, I extend Bayesian estimation with Sequential Monte Carlo sampling to the indeterminacy domain. These findings advance the econometric study of expectations and highlight the macroeconomic relevance of diagnostic beliefs.
JEL classification: C11, C13, C54, C63, E71
Keywords: Diagnostic expectations, DSGE, Identification
@@ 1 Introduction
I begin with a small-scale DSGE framework and show that the diagnostic parameter can be identified globally using the frequency domain methodology of Qu and Tkachenko (2017). This approach introduces a frequency domain expression for the Kullback–Leibler (KL) distance between two DSGE models, which measures the difference between their implied spectra (the Fourier transform of the autocovariance functions). If no two parameter sets generate identical spectrum, that is, if the KL distance is strictly positive, the model is globally identified. Applying this criterion, I find that incorporating DE does not compromise overall model identification but systematically weakens the identification of shock variances. By contrast, key structural parameters remain well identified. This pattern is consistent with the theoretical structure of DE: the diagnostic parameter alters only the shock impact coefficients while leaving autoregressive dynamics unchanged, thereby reducing the sensitivity of the model implied spectrum to changes in shock variances.
@@
Related literature.   This paper relates to two strands of literature: the literature on DE and the broader research on the identification of structural parameters in macroeconomic models. The first strand relates to the growing body of research on DE. Bordalo et al. (2018) (BGS) introduce DE to explain several features of credit cycles, demonstrating how agents’ psychological tendency to overweight representative future outcomes amplifies economic fluctuations. Bordalo et al. (2020) explore DE with dispersed information, finding that agents overreact to private signals while underreacting to consensus forecasts. Bordalo et al. (2021) incorporate price learning and speculative behavior, accounting for the underreaction-overshooting-crash pattern in asset price bubbles, while Bordalo et al. (2021) demonstrate that embedding DE helps account for financial reversals in business cycle models. Guo et al. (2023) examine DE under incomplete information and document excess consumption sensitivity consistent with survey evidence. More recently, Bianchi et al. (2024a) show that a DE-based RBC model better replicates boom-bust cycles compared to its rational expectations counterpart. Na and Yoo (2025) apply this approach to study countercyclical external balances in emerging markets. L’Huillier et al. (2024) incoporate DE into a New Keynesian framework, providing a foundation for my identification analysis. While recent contributions propose smooth DE that distortion vary with uncertainty (Bianchi et al., 2024b), this paper employs the standard BGS framework, which remains the theoretical foundation for most research in this area.
@@ 3.2.4 Global identification
Under DE, the shock variance parameters $\sigma_{a}$ and $\sigma_{g}$ exhibit the weakest identification strength, followed by $\phi_{\pi}$. As shown in L’Huillier et al. (2024)’s online appendix [^9] See also the analytical solution to the modified model in Appendix section A1. A small-scale DSGE model: Analytical solution., the DE parameter $\theta$ influences the policy functions solely through the shock terms. Consequently, introducing $\theta$ effectively redistributes the explanatory power of shocks on endogenous variables, which reduces the curvature of the model implied spectrum to changes in shock variances. Intuitively, since $\theta$ modulates how shocks propagate into the system, variations in the underlying shock variances produce smaller changes in the observable spectral properties. This reduction in spectral sensitivity translates into weaker identification for $\sigma_{a}$ and $\sigma_{g}$. Despite this, the model remains globally identifiable under DE, with the DE parameter $\theta$ itself demonstrating relatively strong identification.
@@ 3.3 Identification under indeterminacy
Notes: The inverse Gamma prior is given by $p(\sigma\mid\iota,s)\propto\sigma^{-\iota-1}e^{-\frac{s^{2}}{2\sigma^{2}}}$, where $\iota=4$ and $s=0.75$. Posterior results were estimated using a modified MATLAB code from Herbst and Schorfheide (2016). All settings are consistent with the determinacy settings in Dynare: number of particles ($N$) = 3,000, number of stages ($N_{\phi}$) = 200, $\lambda=2$, initial scaling parameter = 0.5, and initial acceptance rate = 0.25. Mutation block = 1. The modifications extend the algorithm to cover the indeterminacy region, manage random seeds deterministically across parallel workers to ensure fully reproducible results, and implement numerically stable weight updating via the log-sum-exp transformation to prevent overflow. The estimation results are robust to increasing the number of stages or particles, as well as to employing diffuse priors for the projection coefficients.
@@ 4.1 Global identification
To estimate the model, L’Huillier et al. (2024) use 12 observables: output growth, consumption growth, investment growth, inflation, the interest rate, wage growth, labor growth, and one-period-ahead forecasts of output growth, consumption growth, investment growth, inflation, and the interest rate. However, as noted by Qu and Tkachenko (2023), global identification analysis requires the spectral density to be non-singular. Following Qu and Tkachenko (2023), I therefore work with log deviations from steady states of output, consumption, investment, wage, and labor, rather than their growth rates. Together with inflation and the interest rate, this yields 7 measurement variables in total. For completeness, Appendix A2. A medium scale DSGE contains a detailed discussion of the modified information friction block as well as the full set of model equations.
Using the same approach as in the small-scale DSGE analysis, I take the posterior mean from Bayesian estimation as the benchmark parameter vector, denoted $\gamma_{0}^{med}$. Since the information-friction block is modified, I re-estimate the medium-scale model using standard MCMC sampling. [^11] I adopt the same priors as in L’Huillier et al. (2024). The resulting posterior distribution matches closely that reported in Table 1 of L’Huillier et al. (2024), with the only notable difference being the variance of the noise shock. This is an expected outcome since the signal now relates to the growth rate rather than the level of permanent productivity. The full posterior results are reported in Table 14.
@@ 4.2 Observable Equivalence and the Role of Frictions
First, I examine observable equivalence between the DE and RE specifications. The KL distance is 0.99 for $T=80$ and 1 for $T=150$, indicating that no observable equivalence exists between DE and RE, even with a rich set of frictions. Among the parameters that shift most significantly, the medium-scale RE model requires weaker habit formation, more flexible prices and wages, more volatile markup shocks, and a more dovish monetary policy response to inflation in order to approximate the spectrum of the DE model. In addition, the inverse Frisch elasticity decreases substantially (from 3.71 to 2.41), implying a much more elastic labor supply under RE. These adjustments are intuitive. To replicate the stronger demand responses induced by DE consumers’ extrapolative expectations, RE consumers must rely less on past consumption habits. Since the RE Phillips curves lack extrapolative dynamics, prices must adjust more flexibly than under DE to reproduce observed behavior. Likewise, markup shocks become more volatile in the RE case because, without endogenous amplification from expectations, the model depends more heavily on exogenous shock variances. The sharp increase in labor supply elasticity reflects the RE model’s need for structural amplification to replicate the employment volatility that DE generates through behavioral amplification. Since RE agents do not overreact to shocks, workers must be far more responsive to wage changes to achieve comparable employment fluctuations. Finally, as argued by L’Huillier et al. (2024), when the intertemporal elasticity of substitution equals one, a hawkish monetary policy rule dampens the propagation of noise shocks in RE models. Hence, to mimic DE dynamics, the RE model requires an unrealistically dovish policy rule.
@@
To further study the role of information frictions, I fix $\sigma_{s}$ at alternative levels and allow the remaining parameters of the DE model to adjust in order to fit the benchmark DE spectrum. In particular, I compare a low friction case ($\sigma_{s}=0.1$) with a high friction case ($\sigma_{s}=1$). This exercise highlights how informational frictions reshape both the structural parameters and the degree of DE distortion required to replicate the benchmark dynamics. For the low frictionless case, no observable equivalence is detected: the KL distance is 0.27 for $T=80$ and 0.41 for $T=150$. The parameters remain close to those of the benchmark DE model. In particular, the DE distortion parameter decreases only slightly, from 0.73 to 0.69. The more notable changes occur in investment-related parameters. When DE agents can observe productivity growth more clearly, investment decisions respond less to spurious signals. As a result, the model requires less volatility from marginal efficiency shocks. For the high friction case, no observable equivalence is detected either: the KL distance equals 0.87 for $T=80$ and 0.98 for $T=150$. With severe information frictions, the underreaction channel becomes stronger, which in turn raises the required degree of diagnostic distortion. In this setting, the model relies more heavily on volatile investment specific shocks to generate sufficient fluctuations, while at the same time making capital utilization efficiency more costly to adjust. These adjustments reflect the economy’s need for stronger structural amplification when informational noise is large.
@@
$\gamma_{0}^{med}$ $\theta=0$ $\sigma_{s}=0.1$ $\sigma_{s}=1$ $\kappa_{p}=0.1$ $\kappa_{w}=0.1$ $h=0.1$
KL 0.1137 0.0066 0.0482 0.0746 0.1061 0.6570
$T=80$ 0.9934 0.2730 0.8685 0.9607 0.9908 1.0000
$T=150$ 1.0000 0.4092 0.9808 0.9982 0.9999 1.0000
$\theta$ 0.72 0.00 0.69 0.91 0.35 0.46 0.10
@@
$\frac{\chi^{\prime\prime}(1)}{\chi^{\prime}(1)}$ 5.09 4.93 4.49 6.60 4.99 3.97 4.98
@@
In sum, these experiments show that DE generates dynamics distinct from RE. While other frictions can partially substitute for DE by shifting extrapolation to other channels, none can fully replicate the benchmark DE dynamics. This suggests that DE captures a distinct endogenous amplification mechanism that cannot be mimicked by conventional frictions alone, underscoring its importance in explaining macroeconomic fluctuations.